In a two-tailed test of a hypothesis using the classical method, the critical value is
denoted by
A) ??/2 B) ?? C) ?/2 D) ?
Answer #1

SOLUTION:

In a two-tailed test of a hypothesis using the classical method, the critical value is

Z/2 CORRECT OPTION (A)

CONSIDER ONE EXAMPLE

NOTE:

At 99% confidence level the z is ,

  = 1 - 99% = 1 - 0.99 = 0.01

/ 2 = 0.01 / 2 = 0.005

Z/2 = Z0.005 = 2.576 CRITICAL VALUE